# Privacy-aware Berrut Approximated Coded Computing for Federated Learning

### v1

@@ V-B Matrix multiplication with privacy
$u(z_{j})$ $=\begin{pmatrix}u_{0}(z_{j})&u_{1}(z_{j})&\ldots&u_{K-1}(z_{j})\end{pmatrix}^{T}$ (9)
$=\begin{pmatrix}q_{0}(z_{j})X_{0}+q_{0}(z_{j})q_{K}(z_{j})R_{0}\\ q_{1}(z_{j})X_{1}+q_{1}(z_{j})q_{K+1}(z_{j})R_{1}\\ \vdots\\ q_{K-1}(z_{j})X_{K-1}+q_{K-1}(z_{j})q_{K+T-1}(z_{j})R_{T-1}\end{pmatrix}.$
@@ V-C Privacy Analysis
We use $\mathcal{T}=\{j_{1},\dots,j_{T}\}$ as the set of $T$ colluding nodes, and the corresponding matrix of the coded elements of $X_{i}$ as $L_{\mathcal{T},X_{i}}\triangleq\begin{pmatrix}q_{i}(z_{j_{1}})&q_{i}(z_{j_{2}})&\ldots&q_{i}(z_{j_{T}})\end{pmatrix}^{T}$, and
$\tilde{L}_{\mathcal{T},X_{i}}\triangleq\begin{pmatrix}q_{K+i}(z_{j_{1}})q_{i}(z_{j_{1}})&\ldots&q_{K+i+v-1}(z_{j_{1}})q_{i}(z_{j_{1}})\\ q_{K+i}(z_{j_{2}})q_{i}(z_{j_{2}})&\ldots&q_{K+i+v-1}(z_{j_{2}})q_{i}(z_{j_{2}})\\ \vdots&\ddots&\vdots\\ q_{K+i}(z_{j_{T}})q_{i}(z_{j_{T}})&\ldots&q_{K+i+v-1}(z_{j_{T}})q_{i}(z_{j_{T}})\end{pmatrix},$
of size $T\times v$. The resultant mutual information upper bound is therefore
$I_{L}(T,i)\leq\max_{\mathcal{T}:|\mathcal{T}|=T}\log_{2}|I_{T}+\frac{s^{2}T}{\sigma^{2}_{n}}\tilde{\Sigma}_{\mathcal{T}}^{-1}\Sigma_{\mathcal{T}}|,$
where $\tilde{\Sigma}_{\mathcal{T}}=\tilde{L}_{\mathcal{T}}\tilde{L}_{\mathcal{T}}^{\dagger}$ and $\Sigma_{\mathcal{T}}=L_{\mathcal{T}}L_{\mathcal{T}}^{\dagger}$. Like the arguments in Section III-C, this result is an upper bound that measures the privacy leakage of the scheme. In this case, the privacy metric refers to an individual row of $i$ of the input matrix and $c$ colluding nodes, so we define our privacy metric $\imath_{L}$ for a fixed number $T$ of semi-honest nodes as the averaged version of the mutual information
$\imath_{L}=\frac{\sum_{i=0}^{K-1}I_{L}(T,i)}{K}<\epsilon.$
@@ V-D Large Matrix Multiplication
Multiplication of large matrices arises frequently and plays a key role in many applications of machine learning and artificial intelligence. In this Section, we introduce two simple methods for coded distributed private matrix multiplications (also valid for coded computing of other functions) when the input matrices are large in comparison to the number of nodes collaborating in the computation. A first technique for drastically cutting the computational complexity consist of encoding multiple rows of the matrices using the same interpolation point. Such change does not prevent the scheme to still compute a function, but it allows to do it with a considerably lower number of worker nodes compared to the length of the inputs. A second alternative technique is to split the input matrices vertically in blocks. This has the major advantage of reducing substantially the communication and computation overheads, since, as we have explained, arbitrary functions (including matrix multiplications) can be computed distributively over encoded blocks, and then reconstructed completely in the decoding step.
@@ V-D1 Encoding Multiple rows per point
Until now, the Berrut Secret Sharing scheme takes as a basis the rational function (5). A consequence of the fact that this function encodes one row of the matrix in each interpolation point is that the communications costs decrease, since each encoded share always has length $1$. However, this advantage can become a problem in cases where the number of worker nodes is much smaller than the length of the inputs, as there could be some undecodable rows of the final result.
@@
where $\mathbf{X}^{\prime}_{i}=\begin{pmatrix}X^{T}_{ri}&X^{T}_{ri+1}&\dots&X^{T}_{ri+r-1}\end{pmatrix}^{T}$ with $P=K/r$, $\mathbf{R}^{\prime}_{i}=\begin{pmatrix}R^{T}_{ri}&R^{T}_{ri+1}&\dots&R^{T}_{ri+r-1}\end{pmatrix}^{T}$ with $S=T/r$, and
@@
Thanks to this approach, it is possible to guarantee that $P<N$ by adjusting the number of rows per point, $r$. Aside from this change, the rest of the steps in encoding and decoding remain the same. Achieving a similar gain for matrix multiplications is not so direct, because the encoded matrix has exactly the same size as the input matrix. In substitution of the basis function for each coded row used initially for coded private multiplication, i.e. (11) we suggest to use the modified version is
$u^{\prime}_{i,v}(z)=\mathbf{X}^{\prime}_{i}t_{i}(z)+\sum_{j=P+i}^{P+i+V}\mathbf{R}^{\prime}_{j-P}t_{j}(z),$
@@
$u^{\prime}_{i}(z_{j})=\begin{pmatrix}t_{ri}(z_{j})X_{ri}+q_{ri}(z_{j})t_{K+qi}(z_{j})R_{ri}\\ t_{ri}(z_{j})X_{ri+1}+q_{ri}(z_{j})t_{K+qi}(z_{j})R_{ri+1}\\ \vdots\\ t_{ri}(z_{j})X_{ri+q-1}+q_{ri}(z_{j})t_{K+qi}(z_{j})R_{ri+r-1}\end{pmatrix}.$
@@ V-D2 Block partition of matrices
We assume that there are at least $b^{2}$ client nodes in the system, so that each node receives from the master two shares corresponding $\hat{A}_{x}$ and $\hat{B}_{y}$, for all possible combinations of $x$ and $y$. Identifying node $j$ with its pair of indices $(x,y)$ (a specific block), it will calculate the share
$\hat{f}_{(x,y)}(u(z_{j}))=\begin{pmatrix}\frac{u_{hx}^{(\hat{A}_{x})}(z_{j})u_{hy}^{(\hat{B}_{y})}(z_{j})}{q_{hy}(z_{j})}&\ldots&\frac{u_{hx}^{(\hat{A}_{x})}(z_{j})u_{hy+h-1}^{(\hat{B}_{y})}(z_{j})}{q_{hy+h-1}(z_{j})}\\ \frac{u_{hx+1}^{(\hat{A}_{x})}(z_{j})u_{hy}^{(\hat{B}_{y})}(z_{j})}{q_{hy}(z_{j})}&\ldots&\frac{u_{hx+1}^{(\hat{A}_{x})}(z_{j})u_{hy+h-1}^{(\hat{B}_{y})}(z_{j})}{q_{hy+h-1}(z_{j})}\\ \vdots&\ddots&\vdots\\ \frac{u_{hx+h-1}^{(\hat{A}_{x})}(z_{j})u_{hx}^{(\hat{B}_{x})}(z_{j})}{q_{hx}(z_{j})}&\ldots&\frac{u_{hy+h-1}^{(\hat{A}_{y})}(z_{j})u_{hy+h-1}^{(\hat{B}_{y})}(z_{j})}{q_{hy+h-1}(z_{j})}\end{pmatrix},$
and then adds together all the rows of $\hat{f}_{(x,y)}(u(z_{j}))$ into a single value. The master node will use all the results of the block $(x,y)$ to decode $\hat{A}_{x}^{T}\hat{B}_{y}$. Once all the blocks are decoded, the matrix product can be reconstructed. Obviously, both the multiple encoding per point and block matrix multiplication as described in this Section can be jointly used.
@@ V-E1 Sharing phase
the cost is similar to normal BSS (cf. Section IV-B), but the size of the messages grows linearly with $KL$, where $K$ and $L$ denote the rows and columns of the input matrix, respectively. Two variants can be distinguished: (i) Encoding multiple rows in each interpolation point, for which the analysis is identical and the computational cost is unchanged; (ii) Block partitioning: in this case the length of messages transmitted in the sharing phase is reduced, since the size of blocks is linear in $KH$, with $H=L/b$ the number of columns of each block, and in $b$, the total number of vertical blocks in which the input matrix was divided. Note that $H<l$ for $b>1$.
@@ V-E3 Reconstruction phase
Similar to normal BSS.
When multiple rows encode each interpolation point, the complexity of the decoding step now scales with $K/r$, $2n$ and $rL$, so the complexity is $\mathcal{O}(\frac{K}{r}(2n+rL))$. Instead, dividing the inputs in vertical blocks, the complexity of the decoding step now scales with $b$, $K$, $2n$ and $h$, and it is $\mathcal{O}(bK(2n+h))$. Since $K+T$, the total computational complexity is $\mathcal{O}(K(2n+L)+KL+3K+3T+N+C)$. For the communication cost, $N^{2}$ messages are required, each one of a size that scales with $KL$, and $N$ messages with a size proportional to $\ell$, are exchanged at most. Analyzing the two features for dealing with large inputs, we obtain that: (i) with multiple rows per interpolation point: The total computational complexity is $\mathcal{O}(\frac{K}{r}(2n+rL)+KL+2K+K/r+3T+N+C)$. Regarding communication, $N^{2}$ messages with a size that scales with $KL$, and $N$ messages with a size that scales with $rL$, are exchanged at most; (ii) Input divided in vertical blocks: The total computational complexity is $\mathcal{O}(bK(2n+h)+Kh+3K+3T+N+C)$. As for communication, $N^{2}$ messages with size $Kh$, and $N$ messages with size $h$, are exchanged at most.
@@ VI Results
We provide in this Section numerical results obtained after performing two sets of experiments. First, computations of non-linear functions relevant to FL, comprising the developments of PBACC (Sect. III) and BSS (Sect. IV). Specifically, we tested several common activation functions (ReLU, Sigmoid and Swish) and also typical functions used in non-linear aggregation methods (Binary Step, and Median). Secondly, we conducted tests related to matrix multiplications, demonstrating the adaptations of BSS for this setting (Sect. V). Matrix products with normal and sparse matrices have been computed.
The experimental design is as follows. First, we fix a maximum privacy leakage $\epsilon$ in order to obtain the values of the other parameters to ensure that privacy level is achieve in presence of a semi-honest minority of colluding nodes ($25$% of total). These values are summarized in Table I. After that, we measure the numerical precision in presence of stragglers with the aim of comparing this precision to the one obtained with the original BSS scheme (without any form of privacy). Our goal is calculate the cost (in terms of precision) of adding privacy. Finally, for non-linear computations we also compare BSS with basic differential privacy: an alternative technique to compute non-linear functions. It is worth mentioning that, since we work in a multi-input configuration, the actual computed result $r()$ will be the addition of the function values over each individual input, so $r=\sum_{i=0}^{N-1}f(X^{(i)})$.
TABLE I: Parameters for the experimental settings.
@@
Working under a very pessimistic scenario, where all the encoded private information would be inferred if all the workers collude, the theoretical maximum privacy leakage is $\imath_{L}=H(X)=\log(2s)\approx 14.28$ bits. With the parameters listed in Table I, if a minority of $50$ nodes collude, the maximum information that could be obtained is approximately $0.197$ bits, which means that, at least, $98.62\%$ of information remains secure. For FL settings this is a good security level, but the parameters in Table I can be adjusted accordingly if a tighter privacy is desired.

### v2

@@ V-B Matrix multiplication with privacy
$u(z_{j})$ $=\begin{pmatrix}u_{0}(z_{j})&u_{1}(z_{j})&\ldots&u_{K-1}(z_{j})\end{pmatrix}^{T}=\begin{pmatrix}q_{0}(z_{j})X_{0}+q_{0}(z_{j})q_{K}(z_{j})R_{0}\\ q_{1}(z_{j})X_{1}+q_{1}(z_{j})q_{K+1}(z_{j})R_{1}\\ \vdots\\ q_{K-1}(z_{j})X_{K-1}+q_{K-1}(z_{j})q_{K+T-1}(z_{j})R_{T-1}\end{pmatrix}.$ (9)
@@ V-C Privacy Analysis
We use $\mathcal{C}=\{j_{1},\dots,j_{c}\}$ as the set of $c$ colluding nodes, and the corresponding matrix of the coded elements of $X_{i}$ as $L_{\mathcal{C},X_{i}}\triangleq\begin{pmatrix}q_{i}(z_{j_{1}})&q_{i}(z_{j_{2}})&\ldots&q_{i}(z_{j_{c}})\end{pmatrix}^{T}$, and
$\tilde{L}_{\mathcal{C},X_{i}}\triangleq\begin{pmatrix}q_{K+i}(z_{j_{1}})q_{i}(z_{j_{1}})&\ldots&q_{K+i+v-1}(z_{j_{1}})q_{i}(z_{j_{1}})\\ q_{K+i}(z_{j_{2}})q_{i}(z_{j_{2}})&\ldots&q_{K+i+v-1}(z_{j_{2}})q_{i}(z_{j_{2}})\\ \vdots&\ddots&\vdots\\ q_{K+i}(z_{j_{c}})q_{i}(z_{j_{c}})&\ldots&q_{K+i+v-1}(z_{j_{c}})q_{i}(z_{j_{c}})\end{pmatrix},$
of size $c\times v$. The resultant mutual information upper bound is therefore
$I_{L}(c,i)\leq\max_{\mathcal{C}:|\mathcal{C}|=c}\log_{2}|I_{c}+\frac{s^{2}T}{\sigma^{2}_{n}}\tilde{\Sigma}_{\mathcal{C}}^{-1}\Sigma_{\mathcal{C}}|,$
where $\tilde{\Sigma}_{\mathcal{C}}=\tilde{L}_{\mathcal{C}}\tilde{L}_{\mathcal{C}}^{\dagger}$ and $\Sigma_{\mathcal{C}}=L_{\mathcal{C}}L_{\mathcal{C}}^{\dagger}$. Like the arguments in Section III-C, this result is an upper bound that measures the privacy leakage of the scheme. In this case, the privacy metric refers to an individual row of $i$ of the input matrix and $c$ colluding nodes, so we define our privacy metric $\imath_{L}$ as the averaged version of the mutual information
$\imath_{L}=\frac{\sum_{i=0}^{K-1}I_{L}(c,i)}{K}<\epsilon.$
@@ V-D Large Matrix Multiplication
Multiplication of large matrices arises frequently and plays a key role in many applications of machine learning and artificial intelligence. In this Section, we introduce two simple methods for coded distributed private matrix multiplications (also valid for coded computing of other functions) when the input matrices are large in comparison to the number of nodes collaborating in the computation. A first technique for drastically cutting the computational complexity consist of encoding multiple rows of the matrices using the same interpolation point. Such change does not prevent the scheme to still compute a function, and it allows to get better precisions with a considerably lower number of worker nodes compared to the length of the inputs. A second alternative technique is to split the input matrices vertically in blocks. This has the major advantage of reducing substantially the communication and computation overheads, since, as we have explained, arbitrary functions (including matrix multiplications) can be computed distributively over encoded blocks, and then reconstructed completely in the decoding step.
@@ V-D1 Encoding Multiple rows per point
Until now, the PBSS scheme takes as a basis the rational function (5). A consequence of the fact that this function encodes one row of the matrix in each interpolation point is that the communications costs decrease, since each encoded share always has length $1$. However, this advantage can become a problem in cases where the number of worker nodes is much smaller than the length of the inputs, as the final result will present a huge loss on precision.
@@
where
$\mathbf{X}^{\prime}_{i}=\begin{pmatrix}X_{ri}\\ X_{ri+1}\\ \vdots\\ X_{ri+r-1}\end{pmatrix},\mathbf{R}^{\prime}_{i}=\begin{pmatrix}R_{ri}\\ R_{ri+1}\\ \vdots\\ R_{ri+r-1}\end{pmatrix},$
with $P=K/r$, $S=T/r$, and
@@
Thanks to this approach, it is possible to guarantee that $P<N$ by adjusting the number of rows per point, $r$. Aside from this change, the rest of the steps in encoding and decoding remain the same. Achieving a similar gain for matrix multiplications is not so direct, because the encoded matrix has exactly the same size as the input matrix. In substitution of the basis function for each coded row used initially for coded private multiplication, i.e. (11) we suggest to use the modified version
$u^{\prime}_{i}(z)=\mathbf{X}^{\prime}_{i}t_{i}(z)+\sum_{j=P+i}^{P+i+V}\mathbf{R}^{\prime}_{j-P}t_{j}(z),$
@@
$u^{\prime}_{i}(z_{j})=\begin{pmatrix}t_{ri}(z_{j})X_{ri}+t_{ri}(z_{j})t_{K+ri}(z_{j})R_{ri}\\ t_{ri}(z_{j})X_{ri+1}+t_{ri}(z_{j})t_{K+ri}(z_{j})R_{ri+1}\\ \vdots\\ t_{ri}(z_{j})X_{ri+q-1}+t_{ri}(z_{j})t_{K+ri}(z_{j})R_{ri+r-1}\end{pmatrix}.$
@@ V-D2 Block partition of matrices
We assume that there are at least $b^{2}$ client nodes in the system, so that each node receives from the master two shares
$u^{(\hat{A}_{x})}(z_{j})=\begin{pmatrix}u_{hx}^{(\hat{A}_{x})}(z_{j})&u_{hx+1}^{(\hat{A}_{x})}(z_{j})&\dots&u_{hx+h-1}^{(\hat{A}_{x})}(z_{j})\end{pmatrix},$
$u^{(\hat{B}_{y})}(z_{j})=\begin{pmatrix}u_{hy}^{(\hat{B}_{y})}(z_{j})&u_{hy+1}^{(\hat{B}_{y})}(z_{j})&\dots&u_{hy+h-1}^{(\hat{B}_{y})}(z_{j})\end{pmatrix},$
corresponding $\hat{A}_{x}$ and $\hat{B}_{y}$, for all possible combinations of $x$ and $y$. Identifying node $j$ with its pair of indices $(x,y)$ (a specific block), it will calculate the share
$\hat{f}_{(x,y)}\bigl{(}u^{(\hat{A}_{x})}(z_{j}),u^{(\hat{B}_{y})}(z_{j})\bigr{)}=\begin{pmatrix}\frac{u_{hx}^{(\hat{A}_{x})}(z_{j})u_{hy}^{(\hat{B}_{y})}(z_{j})}{q_{hy}(z_{j})}&\ldots&\frac{u_{hx}^{(\hat{A}_{x})}(z_{j})u_{hy+h-1}^{(\hat{B}_{y})}(z_{j})}{q_{hy+h-1}(z_{j})}\\ \frac{u_{hx+1}^{(\hat{A}_{x})}(z_{j})u_{hy}^{(\hat{B}_{y})}(z_{j})}{q_{hy}(z_{j})}&\ldots&\frac{u_{hx+1}^{(\hat{A}_{x})}(z_{j})u_{hy+h-1}^{(\hat{B}_{y})}(z_{j})}{q_{hy+h-1}(z_{j})}\\ \vdots&\ddots&\vdots\\ \frac{u_{hx+h-1}^{(\hat{A}_{x})}(z_{j})u_{hy}^{(\hat{B}_{y})}(z_{j})}{q_{hy}(z_{j})}&\ldots&\frac{u_{hx+h-1}^{(\hat{A}_{x})}(z_{j})u_{hy+h-1}^{(\hat{B}_{y})}(z_{j})}{q_{hy+h-1}(z_{j})}\end{pmatrix},$
and then adds together all the rows of $\hat{f}_{(x,y)}\bigl{(}u^{(\hat{A}_{x})}(z_{j}),u^{(\hat{B}_{y})}(z_{j})\bigr{)}$ into a single value. The master node will use all the results of the block $(x,y)$ to decode $\hat{A}_{x}^{T}\hat{B}_{y}$. Once all the blocks are decoded, the matrix product can be reconstructed. Obviously, both the multiple encoding per point and block matrix multiplication as described in this Section can be jointly used.
@@ V-E1 Sharing phase
the cost is similar to normal PBSS (cf. Section IV-B), but with only two inputs and the size of the messages grows linearly with $KL$, where $K$ and $L$ denote the rows and columns of the input matrix, respectively. Two variants can be distinguished: (i) Encoding multiple rows in each interpolation point, for which the analysis is identical and the computational cost is unchanged; (ii) Block partitioning: in this case the length of messages transmitted in the sharing phase is reduced, since the size of blocks is linear in $KH$, with $H=L/b$ the number of columns of each block, and in $b$, the total number of vertical blocks in which the input matrix was divided. Note that $H<l$ for $b>1$.
@@ V-E3 Reconstruction phase
Similar to normal PBSS.
When multiple rows encode each interpolation point, the complexity of the decoding step now scales with $K/r$, $2n$ and $rL$, so the complexity is $\mathcal{O}(\frac{K}{r}(2n+rL))$. Instead, dividing the inputs in vertical blocks, the complexity of the decoding step now scales with $b$, $K$, $2n$ and $h$, and it is $\mathcal{O}(bK(2n+h))$. Therefore, the total computational complexity is $\mathcal{O}(K(2n+L)+KL+3K+3T+N+C)$. For the communication cost, assuming two input matrices, $2N$ messages are required, each one of a size that scales with $KL$, and $N$ messages with a size proportional to $L$, are exchanged at most. Analyzing the two features for dealing with large inputs, we obtain that: (i) with multiple rows per interpolation point: The total computational complexity is $\mathcal{O}(\frac{K}{r}(2n+rL)+KL+2K+K/r+3T+N+C)$. Regarding communication, $2N$ messages with a size that scales with $KL$, and $N$ messages with a size that scales with $rL$, are exchanged at most; (ii) Input divided in vertical blocks: The total computational complexity is $\mathcal{O}(bK(2n+h)+Kh+3K+3T+N+C)$. As for communication, $2N$ messages with size $Kh$, and $N$ messages with size $h$, are exchanged at most.
V-F Comparison with other Secure Matrix Multiplication Schemes
TABLE II: Secure distributed matrix multiplication schemes comparison.
protocol privacy precision partition
SCMPC [34] Information-theoretic Exact Vertical
EPC [36] Information-theoretic Exact Complex
FPSCMM [35] Information-theoretic Exact Complex
PBSS Bounded Leakage Bounded Error Vertical
In [34], they propose a scheme for secure massive matrix operations that provides a similar communication complexity as PBSS in the sharing phase, since they both allow to partition the input matrices in vertical blocks, but a worse result sharing phase, as PBSS compresses the result blocks to $1$ row before sending them. The main strengths of [34] against PBSS is that provides information-theoretical privacy in presence of colluding nodes, and an exact computation of the matrix result. In [36] and [35], they propose more efficient methods of secure matrix multiplication, with the possibility of performing complex partitions, rather than just vertical ones. All of the schemes compared resist stragglers better than PBSS as long as the number of colluding nodes or stragglers in the network is not too high, which could impede the results to be reconstructed. In those cases, PBSS could have an advantage if the precision requirements are not too high, as it always reconstructs the result with a given error. Another aspect that should be highlighted is that all of the compared schemes rely on polynomial interpolation for decoding results, which leads numerical instabilities when using real numbers or when the number of nodes of the network is too high. However, the main strength that PBSS has against all the compared schemes is that it is able to compute both matrix multiplications and non-linear functions (really common operations in machine learning). Having achieved a scheme that can compute both operations means that we can deal with the decentralized training of machine learning models with the same scheme. This means that we only need to do one sharing phase, one computation phase, and one decoding phase, and we could reconstruct combinations of non-linear functions and matrix multiplications, while keeping the privacy of the input data. To the extent of our knowledge, our scheme is the most efficient approach that can deal with the two capabilities combined, which means that if we want to securely distribute the training of some ML model in a more efficient way than using PBSS, we would need at least two different schemes (one for dealing with the secure computation of non-linear functions, and the other for the secure computation of matrix multiplications). This imposes two main issues: (i) the more schemes we have to deal with, the harder it is to integrate them in a ML model, and (ii) the schemes in this field always require, at least, some sharing, computation and decoding phases, that would need to be repeated for both schemes and for each function computed, increasing the communication and computational costs of the whole distributed task.
@@ VI Results
We provide in this Section numerical results obtained after performing two sets of experiments. First, computations of non-linear functions relevant to FL, comprising the developments of PBACC (Sect. III) and PBSS (Sect. IV). Specifically, we tested several common activation functions (ReLU, Sigmoid and Swish) and also typical functions used in non-linear aggregation methods (Binary Step, and Median). Secondly, we conducted tests related to matrix multiplications, demonstrating the adaptations of PBSS for this setting (Sect. V). Matrix products with normal and sparse matrices have been computed.
The experimental design is as follows. First, we fix a maximum privacy leakage $\epsilon$ in order to obtain the values of the other parameters to ensure that theprivacy level is achieved in presence of a semi-honest minority of colluding nodes ($c=50$, which represents $25$% of total). These values are summarized in Table III. After that, we measure the numerical precision in presence of stragglers with the aim of comparing this precision to the one obtained with the original BSS scheme (without any form of privacy). Our goal is to calculate the cost (in terms of precision) of adding privacy. Finally, for non-linear computations we also compare BSS with basic differential privacy: an alternative technique to compute non-linear functions. It is worth mentioning that, since we work in a multi-input configuration, the actual computed result $\hat{r}$ will be the addition of the function values over each individual input, so $\hat{r}=\sum_{i=0}^{N-1}f(X^{(i)})$.
TABLE III: Parameters for the experimental settings.
@@
Working under a very pessimistic scenario, where all the encoded private information would be inferred if all the workers collude, the theoretical maximum privacy leakage is $\imath_{L}=H(X)=\log(2s)\approx 14.28$ bits. With the parameters listed in Table III, if a minority of $50$ nodes collude, the maximum information that could be obtained is approximately $0.197$ bits, which means that, at least, $98.62\%$ of information remains secure. For FL settings this is a good security level, but the parameters in Table III can be adjusted accordingly if a tighter privacy is desired.